OpenStudy (anonymous):

can i get some help with this question please find the center of mass of a triangle with points (5,8) (9,11) and (5,11)

OpenStudy (mathmate):

Is this triangle is of uniform density, or there is a mass at each of the vertices?

OpenStudy (anonymous):

uniform density

OpenStudy (mathmate):

Join any vertex to the mid-point of the opposite side. The centre of mass is located at 2/3 of the distance from the vertex.

OpenStudy (anonymous):

would that be all i need to do for this problem

OpenStudy (mathmate):

Yes, but you will do it using analytic geometry.

OpenStudy (anonymous):

im confused because that's not how I learned it in class

OpenStudy (mathmate):

If you go through all the motions, you will end up with the result equivalent to: G((x1+x2+x3)/3, (y1+y2+y3)/3) where the points P1,P2,P3 have coordinates P1(x1,y1),P2(x2,y2),P3(x3,y3), and G is the centre of gravity.

OpenStudy (mathmate):

How did you learn it in class?

OpenStudy (anonymous):

for these type of problems my professor wanted us to use x bar and y bar

OpenStudy (mathmate):

Can you tell me more?

OpenStudy (mathmate):

Do you use calculus?

OpenStudy (anonymous):

yes. this is a cal 2 problem

OpenStudy (mathmate):

OK, there are different ways to do the same thing. You really are learning Calculus, so you need to do integration. What I gave you were known results.

OpenStudy (mathmate):

Are you familiar with xBar = integral(x dA) / integral (dA) and yBar = integral(y dA) / integral(dA) ?

OpenStudy (anonymous):

xbar = 1/A integral xf(x) dx

OpenStudy (anonymous):

y bar = 1/A integral 1/2 [f(x)]^2 dx

OpenStudy (mathmate):

So your f(x) represents the height of the figure at x, right?

OpenStudy (mathmate):

These formulas reduce the double integral (dA) to a simple integral.

OpenStudy (anonymous):

i believe so

OpenStudy (mathmate):

can you first find the height of the figure in terms of x. We will call the three points A, B and C respectively.

OpenStudy (mathmate):

Have you started by drawing the triangle on paper (or calculator)? A is the bottom vertex, B is the right vertex, and C is the top-left vertex.

OpenStudy (anonymous):

yes i have

OpenStudy (mathmate):

We will proceed as follows: 1. We note that the triangle has two sides parallel to the axes, AC // y-axis, BC // to x-axis. 2. We will find the function f(x) which represents the y-dimension of the triangle for x between 5 and 9 3. We will then proceed to do the integrations you suggested. Is this ok with you?

OpenStudy (anonymous):

yes, this is okay with me.

OpenStudy (mathmate):

Can you now start by finding f(x) representing the y-dimension of the triangle. We know that 5<= x <= 9. and f(5)=11-8=3, f(9)=0 It should be a linear function. Confirm the above information on your figure.

OpenStudy (mathmate):

We know that the slope of f(x) is -3/4, since f(5)=3, f(9)=0. So f(x)=-(3/4)x+27/4 Do you agree?

OpenStudy (mathmate):

Now you can proceed with the integrations used in your class to find xBar and yBar.

OpenStudy (mathmate):

Except that the integral for yBar assumes that the base of the triangle is along the x-axis. You can still use the same formula, but the resulting yBar will be measured downwards from the side CB. You can use my formulae to check your results, namely xBar = (x1+x2+x3)/3 yBar = (y1+y2+y3)/3 Good luck.